A direct current of 5 ampere is superimposed on an alternating current of I = 10sinwt what is the effective value of current??

Grade:10

1 Answers

Ritika Chatterjee
6 years ago
Irms^2dt = 5^2dt + (10sinwt)^2dtIrms^2 = [25dt + to 100sin^2wtdt]/dtIrms^2 = (25t + 50t)/tIrms = 75^1/2 = 5 × 3^1/2